OpenStudy (anonymous):

help me pwease =( Women in the United States continue to be paid less than men overall. Nationally, in 2007, the median weekly earnings of female full-time wage and salary workers totaled $614, while it was $766 for males. In 2007, male full-time wage and salary workers earned _________ % more than female full-time wage and salary workers. In 2007, female full-time wage and salary workers earned _________ % less than male full-time wage and salary workers.

OpenStudy (anonymous):

i have the answe btw i just need to verify if its correct

OpenStudy (anonymous):

cheater show me urs first

OpenStudy (anonymous):

no u wont

OpenStudy (anonymous):

Males earned 766/1380= 55.51% more. Females made 614/1380= 44.49% less.

OpenStudy (anonymous):

because thats the whole and u divide by the part

OpenStudy (anonymous):

wth u didnt even help me u NINCOMPOOP

OpenStudy (anonymous):

In 2007, male full-time wage and salary workers earned 24.76 % more than female full-time wage and salary workers. In 2007, female full-time wage and salary workers earned 19.84 % less than male full-time wage and salary workers.

OpenStudy (anonymous):

why was i wrong jeff can u show ur work please

OpenStudy (anonymous):

hmmmm im confused

OpenStudy (anonymous):

The first question was asking for how many percent does the male salary higher than the female salary. So, you must use the female salary as the base, resulting to this equation. \[\frac{ 766-614 }{ 614 } \] = 24.76% Same goes with the second question, but this time since it was asking for the female compared to male, use the male salary as the base. \[\frac{ 766-614 }{ 766 }\] = 19.84%

OpenStudy (anonymous):

ohhhhhhhhhhhhhh i got it lolol thank u thank u !!

OpenStudy (anonymous):

Are you all understanding why you are getting different numbers than Jeff? @nincompoop effectively solved for "what is the ratio of women's pay to men's pay?" @Jeffrey_Calderon solved for "knowing that men were paid more by 766-614, how much of a % difference is that versus the women's pay?" and "how much of a % difference is 766-614 versus the men's pay?"

OpenStudy (anonymous):

Easy number example: If I earn $15 and you earn 10%, I make 150% of your pay. And you make 66.67% of my pay. However, my pay is 50% higher than yours (15-10)/10 = 5/10 = 50% and yours is 33.33% lower than mine (15-10)/15 = 5/15 = 33.33%

OpenStudy (anonymous):

The first approach in that example is what @nincompoop did, and the 2nd approach is what @Jeffrey_Calderon did.

OpenStudy (anonymous):

So @yomamabf , which approach did you choose for your answer? :)

OpenStudy (anonymous):

the guy was right and hey jake =D

OpenStudy (anonymous):

lolol no im fine but i still need my question answered about the ratio thingy

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (anonymous):

i have NO idea what they're asking here

OpenStudy (anonymous):

The way the question is asked, it isn't completely clear... that's why I asked :) Like my example with $10 and $15 pay, there are two totally legit ways to talk about how is paid more and by how much. One is a straight ratio, like @nincompoop , and the other is the ratio of the difference in pay versus one group or the other, like @Jeffrey_Calderon I think @Jeffrey_Calderon is correct here, because it says "men were paid ___% MORE than women" meaning, find how much more they were paid in regular dollars, then calculate that percentage by dividing by the women's pay. But I admit, it is a little confusing the way it's worded..

OpenStudy (anonymous):

yea i thought that too i really thought adding the numbers than dividing by the numbers made sense =/ guess i was wrong

OpenStudy (anonymous):

Lots of problems with discount percentages or "increase in price" percentages follow this form. Shirt costs $20. On sale, shirt costs $15. How much is the % discount? Answer: % discount = (original price - sale price)/original price = (20-15)/20 =5/20 =25%

OpenStudy (anonymous):

Or, inflation strikes, prices rise... $20 shirt last year NOW costs $25. By what % did the shirt increase in price? % increase = (new price - old price)/old price = (25-20)/20 = 25% increase

OpenStudy (anonymous):

can you help me out with the second question jake?

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (anonymous):

Aha!! this is what @nincompoop is solving!! \[\frac{ women }{ men }\]

OpenStudy (anonymous):

614/766 = 80.2

OpenStudy (anonymous):

ohhhh you're supposed to divide..... durrr lol

OpenStudy (anonymous):

so wait you;re just looking for the difference?

OpenStudy (anonymous):

The point is that your first question and your second question are actually different questions with different answers

OpenStudy (anonymous):

i still dont get what the question is aksing to be honest i've read it like 10 times now

OpenStudy (anonymous):

oh so the first 2 answers i solved?

OpenStudy (anonymous):

it doesn't ask you to compare against the mean of men and women... you could find that, but that's not what is asked.

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (anonymous):

690 is the median

OpenStudy (anonymous):

wait why though?

OpenStudy (anonymous):

88.99% i dont understand why tho

OpenStudy (anonymous):

wow thats really confusing wording

OpenStudy (anonymous):

yea just the last part i really dont see where it says you have to divide by the median from the women's earnings

OpenStudy (swissgirl):

I bet yomama doesnt want help from me lol

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (anonymous):

what about the first part where it mentions the male earnings as well?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

hi jeff there was a second part to the question

OpenStudy (anonymous):

holy cow this wording is drivin me nuts

OpenStudy (anonymous):

what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent?

OpenStudy (anonymous):

First, identify the given, what is the male full time wage, salary worker median weekly earnings and female median weekly earnings

OpenStudy (anonymous):

male 766

OpenStudy (anonymous):

median is 766+614= ___ /2 = 690

OpenStudy (anonymous):

female median weekly earnings i have no clue

OpenStudy (anonymous):

the female median is the female earnings?

OpenStudy (anonymous):

oh shnap ohhh i see lol

OpenStudy (anonymous):

ok so whats the next step?

OpenStudy (anonymous):

so what is the final product ?

OpenStudy (anonymous):

wait i dont understand what the final product is supposed to be....

OpenStudy (anonymous):

are you sure of that men plus women over 2?

OpenStudy (anonymous):

so the final product is?

OpenStudy (anonymous):

so there you go yomamabf. I think everything you need to know is here.

OpenStudy (anonymous):

so i should just write that the median for the 2 earnings is 690?

OpenStudy (anonymous):

yes! definitely!

OpenStudy (anonymous):

oh ok ... lol thanx guys <3

OpenStudy (anonymous):

question closed! :)

OpenStudy (anonymous):

oh noooo nincompoop LMAOOO by final product i meant what do i write as the answer

OpenStudy (anonymous):

mother father! haha. you're going to arrive with 2 answers, it's already written above. answer number 1 women median vs men median = women/men answer number 2 women median vs salary worker median = women/([men+women]/2) got it pretty girl? question closed! :)

OpenStudy (anonymous):

they're not asking for the median at all lol i just read the problem corectly they're asking how much do the men make percentage wise from the women lolol

OpenStudy (anonymous):

This is the 2nd question that @yomamabf posed: what percent of male full-time wage and salary worker median weekly earnings does the female median weekly earnings represent? I don't see anything in that question that asks for a comparison with "median worker" It is still a basic ratio of female / male salaries. But this so much fun, I say we just keep solving this dead horse all night :)

OpenStudy (anonymous):

lololol <33 nincompoop

OpenStudy (anonymous):

ROFLLLLLLLLLLLLLLLLLL

OpenStudy (anonymous):

hahahahhahaa

OpenStudy (anonymous):

isn't it that it was also asking for SALARY WORKER MEDIAN WEEKLY EARNINGS?

OpenStudy (anonymous):

nincompoop read the original problem by the median weekly earnings they're just talking about the 766 and 614 lol

OpenStudy (anonymous):

It is the median (male full time wage and male salary) worker... a single male median figure of $766

OpenStudy (anonymous):

theres no median of 690 involved at all lolololol omg we wasted so much time beating a dead horse

OpenStudy (anonymous):

nice jake! Got it! haha. clause problem. haha

OpenStudy (anonymous):

All those words = distractions... it's just 2 figures for men pay and women pay

OpenStudy (anonymous):

darn words... should leave them to the English classes and let us talk pure numbers1!

OpenStudy (anonymous):

OOOOOOKEEEEY! Question closed! Agree?! :)

OpenStudy (anonymous):

I second the motion

OpenStudy (anonymous):

lololol thanx guys <333333333333